OpenStudy (anonymous):

It costs $1.80 less to travel 100 miles in one car than it did to travel 80 miles in another. The average cost per mile of the former is 6 cents less than the average cost per mile of the latter. Find the cost per mile of each.

OpenStudy (anonymous):

Let x be average cost per mile in 1st car and y in the second \[100x +1.80= 80y\] \[x+0.06 = y\] Solving \[100x+1.80 = 80(x+0.06)=80x+4.8\] \[20x=3\] \[x=0.15\] \[y=0.21\]